# Evolution of Shock Structures and QPOs After Halting BHL Accretion onto Kerr Black Hole

### v1

@@ Abstract
One of the mechanisms responsible for disk formation around the black holes is Bondi-Hoyle-Lyttleton (BHL) accretion. The fact that BHL accretion can be interrupted by various astrophysical phenomena, such as stellar winds or astrophysical jets, makes it crucial to study the behavior of shock cones formed by BHL accretion around black holes once the accretion process is halted. Investigating the new plasma structures that emerge in these scenarios can provide insights into observational results. In this context, a new plasma structure forming around the Kerr black hole has been numerically modeled as a function of the black hole’s spin parameter and the asymptotic velocity of BHL accretion. The numerical analysis revealed that high spin ($a/M=0.9$) and supersonic flow ( $M>1$) are necessary conditions for low-frequency quasi-periodic oscillations (LFQPOs) formation. On the other hand, the fundamental mode of the high-frequency quasi-periodic oscillations (HFQPOs) are found to be independent of both the black hole spin and asymptotic velocity and are instead governed by general relativistic effects. Additionally, the study demonstrated that for $3:2$ and $2:1$ resonance states to form, nonlinear couplings needs to be occurred when the black hole rotates rapidly. These couplings produce harmonic frequencies, providing an explanation for the observed quasi-periodic oscillation (QPO) resonances in black hole binaries. These findings align with precession models and nonlinear resonance models, both of which play a crucial role in QPO generation. Finally, the LFQPOs and HFQPOs obtained from numerical simulations are consistent with the observed QPO frequencies in the microquasars $GRS1915+105$ and $XTEJ1550-564$, as well as in the AGN $REJ1034+396$, which harbors a supermassive black hole at its center.
@@
numerical relativity, Kerr metric, BHL, QPOs
@@ 1 Introduction
LFQPOs, calculated from variations in $X-$ray data, are generally observed in black hole binary systems. That is, they arise as a result of the interaction between the black hole and matter, as well as the structure of the accretion disk formed by the infalling material toward the black hole. These QPOs occur both around stellar-mass black holes and around supermassive black holes at the centers of AGNs. For stellar-mass black holes, LFQPO frequencies typically fall within the 0.1300.1300.1–300.1 – 30 Hz range. Observing these frequencies not only allows us to understand the dynamic structure of the disk near the event horizon but also helps reveal the interaction between matter and black holes in strong gravitational fields 2005ApJ…629..403C ; Ingram2009MNRAS ; 2019NewAR..8501524I . LFQPOs are classified into Type-A, Type-B, and Type-C, based on their frequency, coherence, and spectral properties 1999ApJ…514..939W ; Motta2014MNRAS ; 2016AN….337..398M . Among these, Type-C QPOs are typically associated with Lense-Thirring precession in the accretion disk and are commonly observed in the hard and intermediate states of $X-$ray binary systems. Lense-Thirring precession is entirely related to the curvature of spacetime induced by the black hole’s spin Ingram2009MNRAS . Due to this characteristic, LFQPOs provide important insight into general relativistic effects 1998ApJ…492L..59S . LFQPOs have been observed in well-known black hole sources such as $GRS1915+105$, $GROJ1655-40$, $XTEJ1550-564$, and $H1743-322$ Remillard2006ARA&A ; BelloniMNRAS2012 . Furthermore, QPO-like signals have been detected in the AGN $REJ1034+396$, where a QPO frequency in the mHz range has been identified 2011MNRAS.417..250M . Studying LFQPOs is particularly important because they provide information about strong gravitational fields and the regions near black holes. However, their exact physical origin remains uncertain 1999A&A…349.1003T ; 2006ApJ…637L.113S . Therefore, numerical studies of LFQPOs, investigating their physical causes and comparing the results with observations, are essential to advance the scientific literature.
The precessional motions and sudden changes in plasma structures around the stellar-mass and the supermassive black holes, caused by different physical processes, lead to the formation of HFQPOs. These QPOs are observed in both binary systems and Active Galactic Nuclei (AGNs). The oscillation frequencies of these QPOs typically range between 404504045040–45040 – 450 Hz for stellar-mass black holes, whereas for supermassive black holes, they vary from mHz to $\mu$Hz, depending on the black hole mass. It is well established that these QPOs originate from regions close to the event horizon, where strong gravitational fields dominate Remillard2006ARA&A . The detection or observation of QPOs not only helps in understanding the physical properties of black holes but also provides insights into the behavior of matter in strong gravitational fields and the dynamic structure of plasma. Observational studies show that HFQPOs often appear in resonance ratios such as $3:2$. Examples include $GROJ1655-40$ ($300$, $450$ Hz), $XTEJ1550-564$ ($184$, $276$ Hz), and $GRS1915+105$ ($41$, $67$ Hz) Remillard1999ApJ ; Strohmayer2001ApJ ; Misra2004 ; 2012MNRAS.427..595M ; Motta2014MNRAS ; Varniere_2018 ; Liu_2021 ; Majumder_2022 ; Motta2023 . Theoretical explanations suggest that these resonance states arise from different oscillation modes, epicyclic motion, or relativistic precession effects Abramowicz2001A&A ; Ingram2019 . Despite their astrophysical significance, HFQPOs have been observed in only a few sources, and even in those cases, they appear intermittently before disappearing. To confirm the accuracy of these frequencies observationally, long-term X-ray observations are required. By analyzing HFQPO data, we can extract crucial information about black hole mass, spin parameters, and accretion mechanisms Zhang_2017 . To improve the precision of these findings, long-duration observations are essential. Therefore, studying HFQPOs should not rely solely on observational efforts; instead, numerical and theoretical studies are also crucial to uncovering their physical nature. Through these combined efforts, we can gain a deeper understanding of the black hole environments and test the strong gravitational fields more effectively.
@@
In this study, we investigate the physical structure of the plasma that forms in the strong gravitational field around the black hole after the shock cone has developed due to BHL accretion and when BHL accretion is later halted due to specific physical phenomena. We examine how the nature of the newly formed plasma changes depending on the black hole’s spin parameter and the asymptotic velocity used in the BHL accretion process. Additionally, we analyze the precession motion of the plasma in the strong gravitational field and the resulting QPO frequencies caused by spacetime curvature. Finally, we compare these QPOs with observational results, theoretical expectations, and previous QPO calculations in the literature, discussing the contribution of our findings to the field.
@@ 2.2 Initial Conditions
Since we are modeling the interaction between the black hole and the surrounding plasma in the two-dimensional equatorial plane, we set $\theta=\frac{\pi}{2}$. In this study, we first simulate the formation of a shock cone in a strong gravitational field due to the interaction between the black hole and infalling matter through BHL accretion. The shock cone reaches a quasi-steady state at approximately $t=2000$M. It is observed that this steady-state condition persists until around $t=6000$M. After this period, BHL accretion is halted, and the subsequent interaction between the black hole and the surrounding matter leads to the formation of a new plasma structure. For rapidly rotating black holes ($a/M=0.9$), the newly formed plasma is believed to reach a stable configuration between $t=40000$M and $t=80000$M. During this interval, we compute the mass accretion rate and perform a power spectrum analysis based on this accretion behavior. In the case of slowly rotating ($a/M=0.5$) and non-rotating ($a/M=0$) black holes, the plasma is assumed to attain stability between $t=26000$M and $t=46000$M, and similar calculations are conducted.
@@
The computational domain consists of $1024$ points in the radial direction and $512$ points in the azimuthal direction. The radial coordinate extends from $r_{min}=2.3M$ to $r_{max}=100M$, while the azimuthal coordinate spans from 00 to $2\pi$. In both this study and my previous works, it has been demonstrated that the resulting QPO oscillations are independent of the resolution. In the radial direction, outflow boundary conditions are applied both near the black hole’s horizon and at the outer boundary. If gas is injected toward the black hole in the upstream region as part of the BHL accretion process, the physical parameters of the injected matter are automatically copied into the ghost zones. The outflow boundary conditions help prevent numerical artifacts from propagating into the region where the disk or plasma is present. In the azimuthal direction, the periodic boundary condition is implemented.
Table 1: The initial physical variables used are as follows: the black hole’s spin parameter $a/M$, the asymptotic velocity of the matter injected from the outer boundary $V_{\infty}/c$, the Mach number $M=\frac{V_{\infty}}{C_{s}}$, and the rest-mass density of the matter injected from the outer boundary $\rho/M^{2}$, respectively. In all initial conditions, the speed of sound is taken as $C_{s}=0.1$. However, the pressure of the infalling matter are calculated based on the speed of sound and the rest-mass density.
@@ 4 Numerical Results
Figure 1: In the case of a rapidly rotating black hole ($a/M=0.9$), color, contour, and vector plots of the rest mass density are given, showing the shock cone formed by BHL accretion (left column) and the final snapshot of the plasma reaching the steady state after BHL accretion is stopped (right column). Each row of screenshots corresponds to a specific value of asymptotic velocity, clearly demonstrating the complete damping of the shock cone into the black hole and the presence of a newly formed physical mechanism. The screenshots are displayed within the range of $[x_{min},y_{min}]=[-70M,-70M]$ and $[x_{max},y_{max}]=[70M,70M]$.
@@
During BHL accretion, the asymptotic velocity is set at $V_{\infty}/c=0.1$, meaning that the inflowing matter is in a sonic regime as it falls toward the black hole. Because of this, the effect of the black hole’s spin parameter on QPO formation is not observed. In other words, even though BHL accretion is stopped, the dominant fundamental mode remained radial epicyclic oscillations. This is because the accreting matter, in the presence of a strong gravitational field, does not experience significant interaction with the black hole which causes Lense-Thirring precession QPO frequency.
@@
Consequently, when Figs.2, 3, 4, and 5 are compared, it is observed that for the Lense-Thirring effect to become apparent and for the QPO frequencies caused by the curvature of spacetime in the strong gravitational field of the black hole to be observed, the matter must exhibit a supersonic behavior around the black hole. The frequencies generated by the Lense-Thirring effect appear as LFQPOs. However, the radial epicyclic frequency has also been observed in both the sonic and supersonic behaviors of the plasma. These frequencies can be detected as HFQPOs by detectors. Thus, for sources $GROJ1655-40$, $GRS1915+105$, $XTEJ1550-564$, and $H1743-322$ wang20242022 ; BelloniMNRAS2012 ; Strohmayer2001ApJ ; Remillard2006ARA&A ; Varniere_2018 ; Majumder_2022 ; Liu_2021 ; Dhaka_2023 , which exhibit both the LFQPO and the HFQPO behavior, the halted BHL accretion mechanism with supersonic flow modeled here can be proposed as a physical mechanism to explain the observed QPOs.
@@
Since the black hole’s spin parameter affects the curvature of spacetime, it also influences the Lense-Thirring frequency. In this section, the effect of the spin parameter on both LFQPOs and HFQPOs has been analyzed. To examine this effect, Fig.6 presents the mass accretion rates occurring at the spin parameter $a/M=0.5$ under the same asymptotic velocity, $V_{\infty}/c=0.2$ and the power spectrum density analysis generated from this accretion rate. In Fig.6, similar to Fig.3, an asymptotic velocity of $V_{\infty}/c=0.2$ is used, but in this case, a moderately spinning black hole with $a/M=0.5$ is considered. As in the case of the moderately spinning black hole in Fig.6, LFQPOs arising due to the Lense-Thirring effect occur at $4$ Hz and $16.8$ Hz, while HFQPOs at $38$ Hz are generated due to the influence of radial epicyclic motion. The QPO frequencies observed in Fig.6, within a certain error margin, exhibit similarities to the peaks in Fig.3.
@@
To reveal the effect of the black hole’s spin parameter on the generated QPOs, we analyze the QPOs formed around the non-rotating black hole. For this purpose, similar to Figs.3 and 6, the mass accretion rate and the power spectrum analysis for the non-rotating black hole at the same asymptotic velocity of $V_{\infty}/c=0.2$ are presented in Fig.7. As seen in Fig.7, no LFQPOs are observed, while only the HFQPO at $36$ Hz is detected. Additionally, since the black hole does not rotate, the high-frequency peaks arising from nonlinear coupling, which are observed in Figs.3, 4, and 5, do not form. The reason for the absence of nonlinear couplings is the lack of a physical mechanism that could excite the fundamental HFQPO mode and generate new frequencies. These physical mechanisms could include shock cones forming around the black hole, strong spiral shock waves, or the warped spacetime caused by the black hole’s rotation 2008MNRAS.386.2297F .
As a result, it has been demonstrated that the black hole’s spin parameter significantly affects the QPO frequencies. If the plasma exhibits supersonic behavior around the rapidly rotating black hole, both LFQPOs and HFQPOs form. However, as the spin parameter decreases, LFQPOs continue to emerge, while the frequencies caused by nonlinear coupling disappear. In the case of a nonrotating black hole model, as expected, both LFQPOs and the frequencies originating from nonlinear coupling completely vanish if there is no physical mechanism to excite the QPO modes.

### v2

@@ Abstract
In this paper, we investigate for the first time the quasi-periodic oscillations (QPOs) that arise as a result of halting the Bondi-Hoyle-Lyttleton (BHL) accretion mechanism around a Kerr black hole. Unlike previous studies, which focused on shock cones or perturbed tori, we show that stopping BHL accretion leads to the formation of a new plasma structure. On this plasma structure, spiral shock waves are observed to develop. These shock waves are found to excite both low-frequency (LFQPOs) and high-frequency QPOs (HFQPOs). Two key results emerge from our numerical simulations. First, LFQPOs arise only when the halted flow is supersonic, identifying for the first time the velocity dependence of their formation. Second, nonlinear couplings that produce $3:2$, $2:1$, and $1:2:3$ resonances appear exclusively for rapidly rotating black holes ($a/M=0.9$), directly linking spin to resonance states. In addition, it is demonstrated that HFQPOs are strongly influenced by the black hole spin. The generated QPOs are observed to be excited through the modes induced by the black hole spacetime curvature. The frequencies obtained from the numerical simulations align with observations of $GRS1915+105$, XTEJ1550564XTEJ1550-564, and $REJ1034+396$, thereby bridging stellar-mass and supermassive black hole systems within a unified framework. Thus, the newly uncovered physical mechanism associated with halted BHL accretion is capable of explaining both the persistence and the intermittency of QPOs across diverse astrophysical environments.
@@
numerical relativity, Kerr metric, halted BHL, QPOs
@@ 1 Introduction
LFQPOs, calculated from variations in $X$-ray data, are generally observed in black hole binary systems. That is, they arise as a result of the interaction between the black hole and matter, as well as the structure of the accretion disk formed by the infalling material toward the black hole. These QPOs occur both around stellar-mass black holes and around supermassive black holes at the centers of AGNs. For stellar-mass black holes, LFQPO frequencies typically fall within the 0.1300.1–30 Hz range. Observing these frequencies not only allows us to understand the dynamic structure of the disk near the event horizon but also helps reveal the interaction between matter and black holes in strong gravitational fields 2005ApJ…629..403C ; Ingram2009MNRAS ; 2019NewAR..8501524I . LFQPOs are classified into Type-A, Type-B, and Type-C, based on their frequency, coherence, and spectral properties 1999ApJ…514..939W ; Motta2014MNRAS ; 2016AN….337..398M . Among these, Type-C QPOs are typically associated with Lense-Thirring precession in the accretion disk and are commonly observed in the hard and intermediate states of $X$-ray binary systems. Lense-Thirring precession is entirely related to the curvature of spacetime induced by the black hole spin Ingram2009MNRAS . Due to this characteristic, LFQPOs provide important insight into general relativistic effects 1998ApJ…492L..59S . LFQPOs have been observed in well-known black hole sources such as $GRS1915+105$, GROJ165540GROJ1655-40, XTEJ1550564XTEJ1550-564, and H1743322H1743-322 Remillard2006ARA&A ; BelloniMNRAS2012 . Furthermore, QPO-like signals have been detected in the AGN $REJ1034+396$, where a QPO frequency in the mHz range has been identified 2011MNRAS.417..250M . Studying LFQPOs is particularly important because they provide information about strong gravitational fields and the regions near black holes. However, their exact physical origin remains uncertain 1999A&A…349.1003T ; 2006ApJ…637L.113S . Therefore, numerical studies of LFQPOs, investigating their physical causes and comparing the results with observations, are essential to advance the scientific literature.
[…]
@@
In this study, we investigate the physical structure of the plasma that forms in the strong gravitational field around the black hole after the shock cone has developed due to BHL accretion and when BHL accretion is later halted due to specific physical phenomena. We examine how the nature of the newly formed plasma changes depending on the black hole spin parameter and the asymptotic velocity used in the BHL accretion process. Additionally, we analyze the precession motion of the plasma in the strong gravitational field and the resulting QPO frequencies caused by spacetime curvature. Finally, we compare these QPOs with observational results, theoretical expectations, and previous QPO calculations in the literature, discussing the contribution of our findings to the field.
@@ 2.2 Initial Conditions
Since we are modeling the interaction between the black hole and the surrounding plasma in the two-dimensional equatorial plane, we set $\theta=\frac{\pi}{2}$. One of the main reasons for choosing 2D simulations is computational efficiency. This allows us to run each model on very long timescales. In this paper, in order to analyze the quasi-periodic oscillations (QPOs) that arise after the BHL mechanism is halted, the code is evolved up to $80000M$. At the same time, by using 2D simulations, we were able to employ much higher resolution, thereby minimizing the impact of numerical errors and spurious oscillations. As demonstrated in 2011MNRAS.417.2899Z , most of the key features of the shock cone produced by the BHL mechanism can already be captured in 2D simulations. On the other hand, as shown in the literature, 3D simulations generally cannot achieve such high resolution due to computational cost and runtime limitations (2022ApJ…933L…9G, ). Of course, since we did not perform 3D simulations here, vertical instabilities, MRI-driven turbulence, and jet collimation cannot be addressed in this study. However, once the 3D GRMHD code we are currently developing is operational, we will have the opportunity to model these effects and compare the physical outcomes.
In this study, we first simulate the formation of a shock cone in a strong gravitational field due to the interaction between the black hole and infalling matter through BHL accretion. The shock cone reaches a quasi-steady state at approximately t=2000t=2000M. It is observed that this steady-state condition persists until around t=6000t=6000M. After this period, BHL accretion is halted, and the subsequent interaction between the black hole and the surrounding matter leads to the formation of a new plasma structure. For rapidly rotating black holes ($a/M=0.9$), the newly formed plasma is believed to reach a stable configuration between t=40000t=40000M and t=80000t=80000M. During this interval, we compute the mass accretion rate and perform a power spectrum analysis based on this accretion behavior. In the case of slowly rotating ($a/M=0.5$) and non-rotating ($a/M=0$) black holes, the plasma is assumed to attain stability between t=26000t=26000M and t=46000t=46000M, and similar calculations are conducted.
@@
We neglect the effects of magnetic fields and perform purely hydrodynamic simulations. The primary reason for this simplification is that, since we focus on the region very close to the black hole ($r<10M$), the gravitational timescale is much shorter than the magnetic timescale, allowing us to reasonably ignore the influence of magnetic fields. Previous studies have shown that magnetic fields affect the instabilities of shock dynamics, the structure of jets formed in strong gravitational fields, and the overall accretion flow dynamics (2012MNRAS.426.3241N, ). Our goal here is to reveal the type of effects produced when BHL accretion is halted within a GRHD framework. Undoubtedly, including magnetic fields in the calculations would introduce additional effects on the mechanisms at play. For example, magnetic pressure and tension could stabilize or destabilize the shock cone, alter the morphology of the spiral shock waves that we observe in our simulations, and change both the detectability and the actual values of the resulting QPO frequencies. Moreover, the magneto-rotational instability (MRI) could induce turbulence in the accretion flow, which in turn would influence the QPOs. Therefore, we believe that our assumptions are reasonable since the numerical simulations presented here provide not only fundamental physical insight but also capture key processes in a strong gravitational field. However, once we complete the GRMHD code we are currently developing, we will be able to assess the impact of magnetic fields on these phenomena and compare the resulting physical outcomes.
The computational domain consists of $1024$ points in the radial direction and $512$ points in the azimuthal direction. The radial coordinate extends from $r_{min}=2.3M$ to $r_{max}=100M$, while the azimuthal coordinate spans from 0 to $2\pi$. In both this study and my previous works, it has been demonstrated that the resulting QPO oscillations are independent of the resolution. In the radial direction, outflow boundary conditions are applied both near the black hole horizon and at the outer boundary. If gas is injected toward the black hole in the upstream region as part of the BHL accretion process, the physical parameters of the injected matter are automatically copied into the ghost zones. The outflow boundary conditions help prevent numerical artifacts from propagating into the region where the disk or plasma is present. In the azimuthal direction, the periodic boundary condition is implemented.
Table 1: The initial physical variables used are as follows: the black hole spin parameter $a/M$, the asymptotic velocity of the matter injected from the outer boundary $V_{\infty}/c$, the Mach number $M=\frac{V_{\infty}}{C_{s}}$, and the rest-mass density of the matter injected from the outer boundary $\rho/M^{2}$, respectively. In all initial conditions, the speed of sound is taken as $C_{s}=0.1$. However, the pressure of the infalling matter are calculated based on the speed of sound and the rest-mass density.
@@ 4 Numerical Results
The QPO frequencies obtained in this study may originate from the fundamental modes already known in the literature (CruzOsorio2023JCAP, ; Donmez2024arXiv240216707D, ), or from their nonlinear couplings (LL1976, ). Nonlinear couplings occur when radial and azimuthal modes of the plasma, generated by spiral shock waves around the black hole, or modes associated with Lense–Thirring precession due to black hole rotation (Ingram2009MNRAS, ), interact to produce new frequencies through sums, differences, or integer multiples of the base frequencies. In the strong gravitational field of a rotating black hole, such couplings naturally arise as a consequence of differential rotation and the interaction between matter and the curved spacetime. For example, the radial epicyclic mode can couple with the azimuthal epicyclic mode, or with the Lense–Thirring precession mode, to generate resonances such as $2:1$ or $3:2$. These resonances are most often observed as HFQPOs. The presence of strong black hole spin enhances frame dragging and shear, which increases the likelihood of mode–mode interactions. As given in the following, this explains why in our results nonlinear couplings appear for rapidly rotating black holes ($a/M=0.9$), but vanish for moderate or zero spin cases. Theoretically, such mode interactions are consistent with the nonlinear resonance interpretation of QPOs (Abramowicz2001A&A, ; Rezzolla2003MNRAS, ). Therefore, as discussed in detail below and confirmed by our numerical calculations, nonlinear coupling modes can be regarded as one of the primary mechanisms responsible for the generation of HFQPOs.
@@
Figure 1: In the case of a rapidly rotating black hole ($a/M=0.9$), color, contour, and vector plots of the rest mass density are given, showing the shock cone formed by BHL accretion (left column) and the final snapshot of the plasma reaching the steady state after BHL accretion is stopped (right column). Each row of screenshots corresponds to a specific value of asymptotic velocity, clearly demonstrating the complete damping of the shock cone into the black hole and the presence of a newly formed physical mechanism. The screenshots are displayed within the range of $x_{min},y_{min}]=[-70M,-70M]$ and $x_{max},y_{max}]=[70M,70M]$.
@@
During BHL accretion, the asymptotic velocity is set at $V_{\infty}/c=0.1$, meaning that the inflowing matter is in a sonic regime as it falls toward the black hole. Because of this, the effect of the black hole spin parameter on QPO formation is not observed. In other words, even though BHL accretion is stopped, the dominant fundamental mode remained radial epicyclic oscillations. This is because the accreting matter, in the presence of a strong gravitational field, does not experience significant interaction with the black hole which causes Lense-Thirring precession QPO frequency.
@@
Consequently, when Figs.2, 3, 4, and 5 are compared, it is observed that for the Lense-Thirring effect to become apparent and for the QPO frequencies caused by the curvature of spacetime in the strong gravitational field of the black hole to be observed, the matter must exhibit a supersonic behavior around the black hole. The frequencies generated by the Lense-Thirring effect appear as LFQPOs. However, the radial epicyclic frequency has also been observed in both the sonic and supersonic behaviors of the plasma. These frequencies can be detected as HFQPOs by detectors. Thus, for sources GROJ165540GROJ1655-40, $GRS1915+105$, XTEJ1550564XTEJ1550-564, and H1743322H1743-322 wang20242022 ; BelloniMNRAS2012 ; Strohmayer2001ApJ ; Remillard2006ARA&A ; Varniere_2018 ; Majumder_2022 ; Liu_2021 ; Dhaka_2023 , which exhibit both the LFQPO and the HFQPO behavior, the halted BHL accretion mechanism with supersonic flow modeled here can be proposed as a physical mechanism to explain the observed QPOs.
@@
Since the black hole spin parameter affects the curvature of spacetime, it also influences the Lense-Thirring frequency. In this section, the effect of the spin parameter on both LFQPOs and HFQPOs has been analyzed. To examine this effect, Fig.6 presents the mass accretion rates occurring at the spin parameter $a/M=0.5$ under the same asymptotic velocity, $V_{\infty}/c=0.2$ and the power spectrum density analysis generated from this accretion rate. In Fig.6, similar to Fig.3, an asymptotic velocity of $V_{\infty}/c=0.2$ is used, but in this case, a moderately spinning black hole with $a/M=0.5$ is considered. As in the case of the moderately spinning black hole in Fig.6, LFQPOs arising due to the Lense-Thirring effect occur at $4$ Hz and $16.8$ Hz, while HFQPOs at $38$ Hz are generated due to the influence of radial epicyclic motion. The QPO frequencies observed in Fig.6, within a certain error margin, exhibit similarities to the peaks in Fig.3.
@@
To reveal the effect of the black hole spin parameter on the generated QPOs, we analyze the QPOs formed around the non-rotating black hole. For this purpose, similar to Figs.3 and 6, the mass accretion rate and the power spectrum analysis for the non-rotating black hole at the same asymptotic velocity of $V_{\infty}/c=0.2$ are presented in Fig.7. As seen in Fig.7, no LFQPOs are observed, while only the HFQPO at $36$ Hz is detected. Additionally, since the black hole does not rotate, the high-frequency peaks arising from nonlinear coupling, which are observed in Figs.3, 4, and 5, do not form. The reason for the absence of nonlinear couplings is the lack of a physical mechanism that could excite the fundamental HFQPO mode and generate new frequencies. These physical mechanisms could include shock cones forming around the black hole, strong spiral shock waves, or the warped spacetime caused by the black hole rotation 2008MNRAS.386.2297F .
As a result, it has been demonstrated that the black hole spin parameter significantly affects the QPO frequencies. If the plasma exhibits supersonic behavior around the rapidly rotating black hole, both LFQPOs and HFQPOs form. However, as the spin parameter decreases, LFQPOs continue to emerge, while the frequencies caused by nonlinear coupling disappear. In the case of a nonrotating black hole model, as expected, both LFQPOs and the frequencies originating from nonlinear coupling completely vanish if there is no physical mechanism to excite the QPO modes.